OpenStudy (anonymous):

Find the maximum value of the given function. y=sin^3 (x) + cos^3(x)-3sinxcosx. Let t=sinx+cosx. Express y in terms of t 1) find the range of values of t

OpenStudy (primeralph):

do you know calculus?

OpenStudy (anonymous):

dear @agent0smith.. need your help:)

OpenStudy (anonymous):

@primeralph.. not really.. my i know y?

OpenStudy (primeralph):

easier when you know calculus.

OpenStudy (anonymous):

try me:)

OpenStudy (anonymous):

if you knew how to take derivatives this problem would probably be easier because you would take a derivative and find critical points which is where the slope of the original function is = to 0 then you could test values to the left and right of the critical points to judge whether or not they were local maximums or minimums

OpenStudy (anonymous):

f'(x)=3sin(x)^(2)*cos(x)-3cos(x)^(2)sin(x)-3[-sin(x)^(2)+cos(x)cos(x)] someone please check that

OpenStudy (anonymous):

let a=cosx and b=sinx now its like a^3 + b^3 - 3ab = (a+b)(a^2 + b^2 - ab) -3ab now since cos^2x + sin^2x = 1 we can have (a+b)(1-ab) - 3ab

OpenStudy (anonymous):

Dear all, kindly refer to the attached file for your perusal. need your help.

OpenStudy (anonymous):

actually, since: t = sinx + cosx, using sum to product formula, I got \(\large t= 2\sqrt2 sin(x+\frac{\pi}{4})\)

OpenStudy (anonymous):

which part is unclear?

OpenStudy (anonymous):

@mukushla.. have you refer to attached file? i dont know how to get...\[ since t=\sqrt{2} \sin (x+\pi/4)\]

OpenStudy (anonymous):

\[y = \sin ^{3} x + \cos ^{3} x -3sinxcosx = (sinx + cosx)^{3} - 3cosxsinx(cosx+sinx) -3sinxcosx\]

OpenStudy (anonymous):

\[t^{3}-3cosxsinx(1+t) \]

OpenStudy (anonymous):

and also t=sinx+cosx = \[\sqrt{2}(1/\sqrt{2}cosx + 1/\sqrt{2}sinx)\] = \[\sqrt{2}\sin(x+45)\] and hence the values of x+45 is between -180 and +180 x is between -225 and +135 degrees